OpenStudy (anonymous):

What is the slope of any line parallel to the line formed by the equation -45x + 9y = 36

OpenStudy (anonymous):

Make it into standard form for the equation of a line y=mx+b. M is your slope. Any line parallel to this line will have the same slope. So, just find M.

OpenStudy (anonymous):

\[9y = 45x + 36\] \[y = \frac{45}{9}x + \frac{36}{9}\] \[y = 5x +4 \]

OpenStudy (anonymous):

m = 5

OpenStudy (anonymous):

What is the slope of any line parallel to the line formed by the equation -24x – 6y = 12,,i got m=-4x correct

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

m is just -4 though

OpenStudy (anonymous):

-4 slope, negative four slope

OpenStudy (anonymous):

What is the slope of any line perpendicular to the line formed by the equation -7x + 63y = 9,,,,m=-9 correct

OpenStudy (anonymous):

ill medal you man don't worry

OpenStudy (anonymous):

Perpendicular will just be the opposite sign for slope. Slope for that equation is \[\frac{7}{63}\], so perpendicular would be \[-\frac{7}{63}\]

OpenStudy (anonymous):

no m=

OpenStudy (anonymous):

63y = 7x + 9 y = 7/63 x + 9/63

OpenStudy (anonymous):

reduced, that's 1/9 = m, so perpendicular = -1/9 slope

OpenStudy (anonymous):

so how would the answer?? look its cuzz it says to type it in

OpenStudy (anonymous):

-1/9

OpenStudy (anonymous):

\[-\frac{1}{9}\]

OpenStudy (anonymous):

Find the equation of the line that is perpendicular to the line formed by the equation 3x + 2y = -14 and passes through the point (-48, 3).

OpenStudy (anonymous):

post new questions

OpenStudy (anonymous):

close this